The transformation T : z "→" w in the complex plane is defined by w = (az + b)/( z + c), where a, b, c ∈ R. Given that w = 3i when z = −3i, and w = 10 − 4i when z = 1 + 4i, find the values of a, b and c. (a) Show that the points for which z is transformed to z lie on a circle and give the centre and radius of this circle. (b) Show that the line through the point z = 4 and perpendicular to the real axis is invariant under T.
1
Expert's answer
2020-06-02T20:03:19-0400

After substitution of "w = 3i,\, z=-3i" we receive:

"3i=\frac{-3ai+b}{c-3i}\Longrightarrow 3c\,i+9=b-3ai" .

Since "a,\,b,\,c" are real, we get "b=9,\,a+c=0" .

Setting "w=10-4i,\,z=1+4i" we get:

"10-4i=\frac{(1+4i)a+9}{1+4i-a}\Longrightarrow 26+36i-a(10-4i)=a(1+4i)+9"

The latter implies that "36\,i=0" .

Thus, there are no coefficients "a,b,c" satisfying the problem.


(a) We set "w=z" and receive:

"z=\frac{az+b}{z+c}\Longrightarrow z^2+(c-a)z-b=0,\,\,z\neq-c"

Setting "z=x+i\,\,y" we get

"x^2+2xy\,i-y^2+(c-a)(x+i\,\,y)-b=0" .

The latter leads to

"\left\{\begin{array}{l}
 x^2-y^2+(c-a)x-b=0 \\
2xy+(c-a)y=0
\end{array}\right."

Solving the latter, we get

  1. If "D=(a-c)^2+4b>0" then "x=\frac{a-c\pm\sqrt{(a-c)^2+4b}}{2},\quad y=0" ;
  2. If "D=(a-c)^2+4b\leq 0" then "x=\frac12(a-c),\quad y=\pm\frac12\sqrt{-(a-c)^2-4b}."

Now we shall construct a circle for each case.

We set "(x-x_0)^2+(y-y_0)^2=R^2" ,

  1. "D=(a-c)^2+4b>0" . Substitution of points yields the following system

"\left\{\begin{array}{l}
 (x_1-x_0)^2+y_0^2=R^2 \\
(x_2-x_0)^2+y_0^2=R^2
\end{array}\right." with "x_1=\frac{a-c-\sqrt{(a-c)^2+4b}}{2},\,\, x_2=\frac{a-c+\sqrt{(a-c)^2+4b}}{2}"

We subtract the first equation from the second and then solve the obtained equation with respect to "x_0" . We receive "x_0=\frac{x_1+x_2}{2}=a-c." We can then choose an arbitrary "y_0" and "R=\sqrt{\frac14(x_1-x_2)^2+y_0^2}" .

2. "D=(a-c)^2+4b\leq 0" . Substitution of points yields the system:

"\left\{\begin{array}{l}
 (x_1-x_0)^2+(y_1-y_0)^2=R^2 \\
(x_1-x_0)^2+(-y_1-y_0)^2=R^2
\end{array}\right." with "y_1=\frac12\sqrt{-(a-c)^2-4b}."

After substracting of the first equation from the second we receive that "y_1y_0=0" .

If "(a-c)^2+4b=0" we can choose arbitrary "x_0,y_0" and set "R=\sqrt{(x_1-x_0)^2+(y_1-y_0)^2}" . Otherwise, we get "y_0=0" . We choose an arbitrary "x_0" and set "R=\sqrt{(x_1-x_0)^2+y_1^2}" .


(b) All points on the line can be presented as "z=4+y\,i" , "y\in\mathbb{R}". After acting with the transformation we receive

"w=\frac{a(4+y\,i)+b}{4+y\,i+c}" .

We substitute "y=0,\,\,y=1" and receive "w=\frac{4a+b}{4+c}" , "w=\frac{(4a+b)(c+4)+1}{(c+4)^2+1}+\frac{c+4-(4a+b)}{(c+4)^2+1}\,i"

In general case (arbitrary "a,b,c" ), these two points are not of the form "w=4+\tilde{y}\,i" , "{\tilde{y}}\in{\mathbb{R}}" .